a. Write in scientific notation:
(1) 5, 876, 000, 000
(2) 0.00203
b. Write without exponents:
(1) 7.5 x 105 (to the fifth)
(2) 7.8 x 10 -4 (to the negative 4th pwr)

Answer by checkley71(8403)

5,876,000,000=5.876*10^8
0.00203=2.03*10(^-3)
7.5*105^5=910,529,667.5
7.8*10(^-4)=7.8*.0001=.00078